# Supermassive black holes from inflation constrained by dark matter substructure

### v1

@@ IV Results and discussion
Upcoming surveys such as the Rubin Observatory Legacy Survey of Space and Time (LSST) Shipp et al. (2023) can extend these constraints to much smaller scales. In Fig. 5, we showcase examples assuming ${N_{\rm sh}(A,k_{p},n_{b})}/{N_{\rm sh}(A=0)}<2$ for four subhalo mass ranges, $10^{5}$$10^{6}$, $10^{6}$$10^{7}$, $10^{7}$$10^{8}$, and $10^{8}$$10^{9}M_{\odot}$ with light to dark magenta. Here we parametrize the abundance ratio by taking the ratio of the abundance to the case with $A=0$, i.e. without an enhancement in the power spectrum. Using this approach, we highlight that there is strong potential to explore the PBH origin of the SMBH by probing its mass down to $\simeq{\cal O}(10^{5})M_{\odot}$ for $n_{b}=4$ and $\simeq{\cal O}(10^{7})M_{\odot}$ for $n_{b}=8$. SMBH seeds of mass $10^{5}$$10^{6}M_{\odot}$ would naturally explain the observed black holes at high redshifts by JWST. In this approach, we can probe power spectrum amplitudes down to $\simeq{\cal O}(10^{-7})$ for a certain range of scales. The constraints can be comparably strong to those expected with future $\mu$-distortion measurements by PIXIE Ünal et al. (2021), which are shown with dashed lines in the figure. Also for $n_{b}=8$, the DM substructure probe can be advantageous relative to to $\mu$-distortion measurements. We also remark in Fig. 6 the relative strength of the current DM substructure constraints for small $k_{p}$ to future $\mu$-distortion probes with PIXIE.
@@
V Conclusion
@@
In such situations, the halo mass function of subhalos around the Milky Way would also be altered, leading to different numbers of satellite galaxies and smaller dark subhalos. We investigated the region of parameter space which can be constrained from current observations on the number counts of the satellite galaxies and dark subhalo estimates with stellar stream and lensing observations. The DM substructure analysis excludes large regions of amplitude for peak wavenumbers less than ${\cal O}(100)$ Mpc-1, corresponding to PBH seed masses greater than $\simeq{\cal O}(10^{9})M_{\odot}$, which cannot be probed with current CMB $\mu$-distortion measurements.
Potential for future observations to extend the probable range with the scheme of this article is also discussed. We highlight the possibility of probing the PBH origin of SMBHs down to masses of approximately ${\cal O}(10^{5})M_{\odot}$ and power spectrum amplitudes as low as ${\cal O}(10^{-7})$ over certain scales. We have shown that DM substructure bounds are comparably powerful to future CMB $\mu$-distortion measurments for probing the enhanced curvature perturbations. Recent works indicate that the second-order gravitational waves (e.g. Ref. Byrnes et al. (2019)) and Lyman-$\alpha$ data (e.g. Ref. Ragavendra et al. (2024)) could also be important for probing certain models which predicts enhanced curvature perturbations. Those probes may have comparable sensitivity in some regions to our work, however, we remark that the prospects shown here with DM substructures are conservatively obtained. If we include propagation of non-Gaussian effects in our DM substructure analysis, we expect it could be even more constraining.
@@ VI Acknowledgements
SB and MF are supported by the STFC under grant ST/X000753/1. SA and MF are extremely grateful for support via the Royal Society International Exchange project “Probing particle nature of dark matter using small-scale distribution” which made this work possible. The work of SA was also supported by MEXT KAKENHI under grant numbers JP20H05850, JP20H05861, and JP24K07039. The work of NH was partly supported by MEXT KAKENHI Grant Number 20H05852, 22K14035, and MEXT Leading Initiative for Excellent Young Researchers Grant Number 2023L0013. The work of KI was supported by JSPS KAKENHI Grant Number JP20H01894 and JSPS Core-to-Core Program Grant No. JPJSCCA20200002.

### v3

@@ IV Results and discussion
Upcoming surveys such as the Rubin Observatory Legacy Survey of Space and Time (LSST) Shipp et al. (2023) can extend these constraints to much smaller scales. In Fig. 5, we showcase examples assuming ${N_{\rm sh}(A,k_{p},n_{b})}/{N_{\rm sh}(A=0)}<2$ for four subhalo mass ranges, $10^{5}$$10^{6}$, $10^{6}$$10^{7}$, $10^{7}$$10^{8}$, and $10^{8}$$10^{9}M_{\odot}$ with light to dark magenta. Here we parametrize the abundance ratio by taking the ratio of the abundance to the case with $A=0$, i.e. without an enhancement in the power spectrum. Using this approach, we highlight that there is strong potential to explore the PBH origin of the SMBH by probing its mass down to $\simeq{\cal O}(10^{5})M_{\odot}$ for $n_{b}=4$ and $\simeq{\cal O}(10^{7})M_{\odot}$ for $n_{b}=8$. SMBH seeds of mass $10^{5}$$10^{6}M_{\odot}$ would naturally explain the observed black holes at high redshifts by JWST. In this approach, we can probe power spectrum amplitudes down to $\simeq{\cal O}(10^{-8})$ for a certain range of scales. The constraints can be comparably strong to those expected with future $\mu$-distortion measurements by PIXIE Ünal et al. (2021), which are shown with dashed lines in the figure. Also for $n_{b}=8$, the DM substructure probe can be advantageous relative to to $\mu$-distortion measurements. We also remark in Fig. 6 the relative strength of the current DM substructure constraints for small $k_{p}$ to future $\mu$-distortion probes with PIXIE.
V Discussion and conclusions
@@
In such situations, the mass function of subhalos around the Milky Way would also be altered, leading to different numbers of satellite galaxies and smaller dark subhalos. We investigated the region of parameter space which can be constrained from current observations on the number counts of the satellite galaxies and dark subhalo estimates with stellar stream and lensing observations. The DM substructure analysis excludes large regions of amplitude for peak wavenumbers less than ${\cal O}(100)$ Mpc-1, corresponding to PBH seed masses greater than $\simeq{\cal O}(10^{9})M_{\odot}$, which cannot be probed with current CMB $\mu$-distortion measurements.
For models with a peaked power spectrum, subhalos corresponding to the peak mass scale are expected to have significantly higher concentrations. These enhanced densities can influence a variety of astrophysical observables including the ones explored in this work. While our satellite constraints indirectly include such structural information as the models are constrained by the number of satellites with $V_{\rm max}>4\penalty 10000\ \mathrm{km\penalty 10000\ s^{-1}}$, we primarily focused on the subhalo mass function. We acknowledge that a full analysis incorporating the internal density structure of subhalos would provide additional insights (e.g., Esteban et al. (2024)).
Potential for future observations to extend the probable range with the scheme of this article is also discussed. We highlight the possibility of probing the PBH origin of SMBHs down to masses of approximately ${\cal O}(10^{5})M_{\odot}$ and power spectrum amplitudes as low as ${\cal O}(10^{-7})$ over certain scales. We have shown that DM substructure bounds are comparably powerful to future CMB $\mu$-distortion measurements for probing the enhanced curvature perturbations. Recent works indicate that the second-order gravitational waves (e.g. Ref. Byrnes et al. (2019)) and Lyman-$\alpha$ data (e.g. Ref. Ragavendra et al. (2024)) could also be important for probing certain models which predict enhanced curvature perturbations. Those probes may have comparable sensitivity in some regions to our work, however, we remark that the prospects shown here with DM substructures are conservatively obtained. There has recently been an effort to propagate non-Gaussian effects onto the $\mu$-distortion bound in Refs. Sharma et al. (2024); Byrnes et al. (2024). However, the effects found in those works are significant only at wavenumbers $k_{p}\gtrsim 10^{4}$-$10^{5}$ Mpc-1, which is complementary to our work. It should be noted though that the limits derived in those works are not directly comparable to ours because of the different templates used for the power spectrum. PBHs and DM subhalos originate from the same primordial power spectrum, hence both of their respective formations are influenced by the shape of the probability distribution function of these fluctuations. Non-Gaussianity that thickens the tails of the probability distribution function (as studied in this work) where large fluctuations occur, would result in a larger number of regions being dense enough to collapse into DM subhalos. This would further strengthen our bounds if we included such effects in our analysis. We intend to study the effect of non-Gaussianity on DM subhalos in detail in a future work while presenting the most conservative scenario in this one.
Note: While this work was under revision, Ref. Nadler et al. (2025) presented numerical simulations of subhalo abundances for inflationary spectra containing moderate log-normal bumps at wavenumbers $k\simeq 30\penalty 10000\ {\rm Mpc^{-1}}$. They reported no significant suppression of low-mass halos, which is consistent with our predictions (see Fig. 3), where the subhalo mass function remains essentially unmodified for comparable amplitudes and scales. Their results therefore complement our semi-analytical analysis by confirming the expected behavior in the weak power spectrum enhancement regime.
@@ VI Acknowledgements
We would like to thank Antonio Iovino, Gabriele Franciolini, Christian Byrnes and Devanshu Sharma for useful feedback. We are grateful to Masahiro Kawasaki for valuable and insightful discussions. SB and MF are supported by the STFC under grant ST/X000753/1. SA and MF are extremely grateful for support via the Royal Society International Exchange project “Probing particle nature of dark matter using small-scale distribution” which made this work possible. The work of SA was also supported by MEXT KAKENHI under grant numbers JP20H05850, JP20H05861, and JP24K07039. The work of NH was partly supported by MEXT KAKENHI Grant Number 20H05852, 22K14035, and MEXT Leading Initiative for Excellent Young Researchers Grant Number 2023L0013. The work of KI was supported by JSPS KAKENHI Grant Number JP20H01894, JP25K07317, and JSPS Core-to-Core Program Grant No. JPJSCCA20200002.